In an instant lottery, your chances of winning are 0.2. If you play the lottery one time and outcomes are independent, the probability that you do NOT win is:
Answer #1

Here ,in an instant lottery, my chances of winning are 0.2. i.e.

P(chances of winning) = p = 0.2

If I play the lottery one time and outcomes are independent.

Then,

The probability that I do NOT win :

P(Do not win) =1 - p = 1 - 0.2 = 0.8

For further query please comment below.thank you